Circular No. 7520

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD B. Sugerman, S. Lawrence, and A. Crotts, Columbia University, report

on evolution and new structure in the ejecta’s interaction with the inner equatorial ring (ER) of the remnant of SN 1987A: “Hubble Space Telescope STIS F28X50LP imaging and G750M spectroscopy (52′′ × 2′′ slit) were obtained on Oct. 29 and Nov. 3 and compared to earlier STIS observations. Using the notation of Lawrence et al. (2000, Ap.J. 537, 123), the first seven hot spots are reconfirmed, and lightcurves of spots HS 2-104, HS 3-126, HS 4-091 and HS 5-139 (all first detectable between Nov. 1998 and Jan. 1999) reveal that they do not evolve uniformly. HS 2-104 was a factor of 2 fainter than HS 3-126 until Oct. 1999, but it is a factor of 2 brighter in Nov. 2000. HS 5-139 has increased sixfold in brightness between May and Nov. 2000, while HS 4-091 has increased by a factor of only 1.4. Two new spots, detected in imaging and the Hα spectrum, are reported as follows with designation, p.a., radius, and the STIS flux (counts/s) for the ER: HS 10-040, 40^{o}, 0′′.64, 7.8 ± 1.0; HS 11-355, 355^{o}, 0′′.52, 2.0 ± 0.4. HS 10-040 is detectable as early as Oct. 1999, and the ‘faint, asymmetric extension’ of HS 1-029 reported on IAUC 7419 is probably best explained by this distinct interaction site just east of HS 1-029. The radial positions and slow evolution of spots HS 8-064 and HS 9-075 (detected at low levels by Lawrence et al. 2000) appear more consistent with reverse-shock emission, rather than ejecta-ER interaction. Additional brightening (at low levels consistent with known spots at very early times but requiring confirmation) is detected at p.a. 49^{o}, 174^{o}, 248^{o}, and 314^{o}.”

COMETS C/2000 V1 and C/2000 V2 (SOHO) Further to IAUC 7514 and 7485, D. Hammer reports his measurements

for two additional Kreutz sungrazing comets imaged with the SOHO spacecraft; C/2000 V1 was found by M. Boschat, and C/2000 V2 was found by D. Biesecker. Both comets were seen in the C2 coronagraph, but only only C/2000 V1 was also visible with the C3 instrument. The reduced observations and orbital elements by B. G. Marsden are given on MPEC 2000-V35.

Comet  2000 UT  α_{2000}  δ_{2000}
C/2000 V1  Nov. 4.471  14^{h}30.^{m}2  −17^{o}29′
C/2000 V2  5.643  14 42.1  −17 24
2000 November 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green